### v1

# Towards SLO-Optimized LLM Serving via Automatic Inference Engine Tuning

Abstract
A service-level objective (SLO) is a target performance metric of service that cloud vendors aim to ensure. Delivering optimized SLOs can enhance user satisfaction and improve the competitiveness of cloud vendors. As large language models (LLMs) are gaining increasing popularity across various fields, it is of great significance to optimize SLOs for LLM inference services. In this paper, we observe that adjusting the parameters of LLM inference engines can improve service performance, and the optimal parameter configurations of different services are different. Therefore, we propose SCOOT, an automatic performance tuning system to optimize SLOs for each LLM inference service by tuning the parameters of the inference engine. We first propose a generalized formulation of the tuning problem to handle various objectives and constraints between parameters, and SCOOT exploits the Bayesian optimization (BO) technique to resolve the problem via exploration and exploitation. Moreover, SCOOT adopts a random forest to learn hidden constraints during the tuning process to mitigate invalid exploration. To improve the tuning efficiency, SCOOT utilizes the parallel suggestion to accelerate the tuning process. Extensive experiments demonstrate that SCOOT can significantly outperform existing tuning techniques in SLO optimization while greatly improving the tuning efficiency.
Index Terms:
large language model, inference engine, service-level object, performance tuning, machine learning
I Introduction
With the impressive capabilities demonstrated by large language models (LLMs) across various applications [1], cloud vendors such as Alibaba Cloud and AWS have started offering services of LLM deployment and inference [2, 3]. Customers (i.e., application developers and providers) can deploy specified LLMs on their exclusive computing resources. Through the application programming interface (API) provided by the cloud vendor, customers‘ requests can be continuously served by the deployed LLMs.
These LLM inference services run LLM instances with advanced inference engines such as vLLM [4] and TensorRT-LLM [5], which are equipped with cutting-edge technologies, such as continuous batching [6], paged attention [7], and chunked prefill [8]. These techniques can accelerate inference and improve throughput, thus delivering a high service performance to customers. To provide flexibility of use, inference engines expose many parameters, of which the values can be modified to enable a particular technique or to change strategies of request scheduling and memory allocation.
To ensure service performance, customers always agree on service-level objectives (SLOs) with cloud vendors. SLOs are defined as a series of performance metric constraints, such as requiring that the 95th percentile latency of requests be less than 1 second. A tight SLO can gain a reputation for cloud vendors and attract more customers. Therefore, it is significant for cloud vendors to optimize SLOs (e.g., guarantee a lower tail latency) to improve customer satisfaction while enhancing market competitiveness.
Existing LLM deployment solutions usually employ the default parameter configuration or adopt an empirically good configuration of the LLM inference engine for all inference services. However, in this paper, we observe that adjusting the parameters of LLM inference engines has great potential to improve service performance, and the optimal parameter configurations of various services are different. Therefore, we stand in the shoes of cloud vendors and optimize SLOs for each LLM inference service by tuning the parameters of the LLM inference engine.
Numerous performance tuning methods have been proposed and applied across various fields [9], but they all fall short of efficiency and optimality for tuning LLM inference engines. Methods like random sampling and meta-heuristic algorithms, such as Monte Carlo sampling [10] and genetic algorithms [11] lack efficiency as they fail to fully utilize historical information. Besides, heuristic searches often rely on expert knowledge and elaborate pre-profiling to model relationships between parameters and performance. Nevertheless, with the rapid evolution of technologies in LLM inference engines, the parameters’ numbers and ranges are frequently updated, which makes the modeled parameter-performance relationship outdated and the designed heuristic search methods inapplicable. Learning-based approaches such as reinforcement learning (RL) [12] and Bayesian optimization (BO) [13] have also been widely exploited in performance tuning. They can effectively leverage historical information and tune parameters automatically without prior knowledge. However, existing learning-based tuning methods have difficulties tuning LLM inference engines due to the following three unique challenges.
Challenge 1: Customers seek to enhance different performance metrics depending on their applications, such as improving request throughput for the offline recommendation application, reducing request tail latency for the online classification application, and minimizing time-to-first-token (TTFT) and time-per-output-token (TPOT) simultaneously for interactive applications such as chatbot. Challenge 2: The tuning process is complicated by both known constraints that can be provided to the tuner ahead of time and hidden constraints that are initially unknown and must be learned throughout the tuning process. Challenge 3: Tuning methods always require significant time to evaluate different parameter configurations, making the tuning process for a large number of LLM inference services time-prohibitive.
To address the above challenges, we propose SCOOT, a ServiCe-level Objective Oriented performance Tuning system, which automatically tune parameters of LLM inference engines to optimize SLOs for LLM inference services. We first propose a general formulation of the inference engine tuning problem to accommodate various optimization objectives and complex constraints, and SCOOT can resolve the problem with BO, where single-objective Bayesian optimization (SOBO) and multi-objective Bayesian optimization (MOBO) [14, 15, 16] are respectively employed to search optimized parameter configurations in single-objective and multi-objective situations, thus addressing challenge 1. Moreover, SCOOT prunes the search space with known constraints and exploits a random forest to learn hidden constraints during the tuning process, thus mitigating challenge 2. To resolve challenge 3, the parallel suggestion is implemented on top of BO to recommend multiple parameter configurations for evaluation each time, thus fully utilizing idle computing resources to greatly accelerate the tuning process and improve the tuning efficiency.
We conduct extensive experiments with various LLMs and different types and numbers of GPUs under request traces collected from various applications in the production environment of Ant Group. The results show that SCOOT can speed up the tuning process and significantly optimize SLOs, improving the request throughput by up to 68.3%, reducing the request tail latency by up to 40.6%, and reducing the TTFT and TPOT by up to 99.8% and 61.0%, respectively, compared to the default parameter configuration and existing tuning methods. The main contribution of this paper is listed as follows.
  • To the best of our knowledge, this is the first study that introduces performance tuning into the field of LLM serving, and we uncover the significance of tuning LLM inference engines with real-world request traces.
  • We propose a general formulation of the inference engine tuning problem that accommodates various optimization objectives and constraints, and we design SCOOT to solve the problem by intelligently searching optimized parameter configurations with BO.
  • Random forest regression is employed by SCOOT to learn hidden constraints during the tuning process to avoid invalid explorations while the parallel suggestion technique is adopted to significantly improve the tuning efficiency using additional computing resources.
  • Extensive experiments are conducted to confirm the superiority of SCOOT in terms of both the optimality and efficiency for tuning LLM inference engines under various LLMs, computing resources, and request traces.
II Background and Motivation
II-A Research Background
TABLE I: PARAMETERS TO TUNE.
Configuration Parameter Type Range
tensor-parallel Integer [1, #GPUs]
max-num-seqs Integer [64, 8192]
max-num-batched-tokens Integer [64, 8192]
block-size Enumeration {8, 16, 32}
scheduler-delay-factor Float [0, 2]
enable-chunked-prefill Boolean {True, False}
enable-prefix-caching Boolean {True, False}
disable-custom-all-reduce Boolean {True, False}
use-v2-block-manager Boolean {True, False}
II-A1 LLM Inference Engine and Parameters
Inference engines integrate numerous advanced LLM inference techniques such as continuous batching [6], paged attention [7], and chunked prefill [8] and expose many parameters. For vLLM, these parameters include boolean variables such as enable-chunked-prefill that can enable the technique of chunked prefill, integer and float variables such as max-num-seqs and scheduler-delay-factor that can change the request scheduling strategy of continuous batching, and enumeration variables such as block-size that can change the memory allocation policy of paged attention. In this paper, we focus on tuning parameters that do not affect the accuracy of LLMs. Therefore, we do not consider parameters related to model compression such as quantization. Additionally, although speculative decoding has been theoretically proven not to hurt LLM accuracy [17], it still affects the LLMs’ generation results, making it inapplicable for certain applications. Hence, we also do not tune parameters related to speculative decoding. In this paper, we choose vLLM as the inference engine, and the parameters to be tuned are listed in Table I, which constructs a huge search space of billions of configuration points.
II-A2 Service-level Objective and Stress Testing
To ensure service performance, customers establish SLOs with cloud vendors. If SLOs are violated, cloud vendors not only have to compensate customers but also face reputational damage. Additionally, customers also need to make a commitment to the request rate, guaranteeing that the request rate for the inference service will not exceed an agreed threshold. Typically, cloud vendors stress test services with a high request rate (i.e., a request rate that is close to the agreed threshold) and set the performance metrics achieved under the stress test as SLOs. This ensures that SLOs won’t be violated even under heavy workload scenarios. By optimizing the service performance under stress testing, cloud vendors can deliver better SLOs to customers.
TABLE II: VARIOUS SERVICE CHARACTERISTICS.
Service Application GPU Type GPU Number LLM
A SQL A100 2 LLAMA2-13B
B BOT A100 2 LLAMA2-13B
C BOT A10 2 LLAMA2-13B
D BOT A10 4 LLAMA2-13B
E BOT A10 4 LLAMA2-7B
II-B Experimental Observations
Figure 1: Optimal TTFT and TPOT for various services. The TTFT and TPOT shown are relative values compared to those of the default parameter configuration. The lower, the better.
II-B1 Parameter adjustment of inference engine can enhance performance for LLM inference services
We conduct experiments to find the optimal parameter configurations of vLLM with a grid search for five services of various LLMs deployed on different numbers and types of GPUs under request traces of various applications. The service characteristics in the experiments are shown in Table II, where A10 and A100 separately represent the NVIDIA A10 24G GPU and NVIDIA A100 80G GPU, and SQL and BOT represent the request traces of text-to-sql and chatbot applications in Ant Group. Other experimental settings are the same as those in Section V. Fig.1 presents the relative TTFT and TPOT performed by optimal parameter configurations for these services. We can observe that TTFT and TPOT can be reduced by up to 98.9% and 49.9% compared to the default configuration, respectively, which confirms the significance of performance tuning.
II-B2 Different inference services’ optimal parameter configurations are different
We conduct experiments to apply the optimal parameter configuration of a service to other services. Figure 2 presents the experimental results. We can observe that the performance exhibited by the optimal parameter configuration for a service may perform poorly or even cause errors for other services. Hence, there is no one “best practice” configuration that works best in all scenarios, and it is essential to conduct performance tuning for each inference service.
II-C Challenges
II-C1 Various Optimization Objectives
Customers of LLM inference services intend to employ different performance metrics as objectives for SLO optimization according to their specific application requirements. For non-interactive applications, periodically invoked offline applications such as offline recommendation, prioritize improving request throughput, while online applications such as online classification, aim to reduce request tail latency. Interactive applications such as chatbot continuously display new tokens generated by the LLM to application users, and hence they tend to simultaneously minimize two token-level metrics, time-to-first-token (TTFT), and time-per-output-token (TPOT). TTFT is the time from the user sending a request to receiving the first token, and TPOT is the average time interval between tokens. By minimizing TTFT and TPOT, the output progress can be more fluent, thereby enhancing the user experience. Therefore, optimizing SLOs for non-interactive applications often poses a single-objective optimization problem, whereas optimizing SLOs for interactive applications is a multi-objective problem.
II-C2 Complex Known and Hidden Constraints
Some parameters of LLM engines depend on the settings of other parameters, thus causing constraints on the search space. For example, for vLLM, max-num-batched-tokens must be greater than or equal to max-num-seqs. If these constraints are violated, the inference engine can not be started. We refer to these as known constraints, which can be provided to the tuner ahead of time. In addition, given a specific service, certain parameter combinations can lead to inference engine crashes during the stress testing. In this paper, these infeasible parameter combinations are referred to as hidden constraints. Different services have different hidden constraints that are initially unknown and must be learned throughout the tuning process. Specifically, for some services, vLLM often crushes due to a timeout error during the stress testing when scheduler-delay-factor is set to a large value.
(a) Relative time-to-first-token (TTFT) compared to default configuration. The lower, the better.
(b) Relative time-per-output-token (TPOT) compared to default configuration. The lower, the better.
Figure 2: TTFT and TPOT for applying optimal parameter configurations of different services to other services. Conf. $X$ represents the optimal parameter configuration for the service $X\in\{A,B,C,D,E\}$. The red lines (1.0×) in Sub-figures (a) and (b) indicate the TTFT and TPOT under the default parameter configuration of the inference engine, respectively.
II-C3 High Evaluation Overhead
Tuning methods always necessitate evaluating various parameter configurations. For the inference engine tuning, each evaluation (i.e., stress testing) can take several minutes. For example, it takes about 6 minutes to stress test a service with the LLAMA2-7B LLLM [18] deployed on 2 A10 GPUs using 1,000 requests of the BOT application. For more requests and larger LLMs, the time required for stress testing will become longer. Hence, even with only 30 evaluations, the time spent tuning a single service can reach several hours, making the time required for tuning a multitude of newly deployed services prohibitive.
II-D Limitation of Existing Solutions
II-D1 Random Sampling and Meta-Heuristics
Random sampling randomly samples points from the search space to evaluate while meta-heuristics such as genetic algorithms [11] and particle swarm optimization [15] first sample several points, and then optimize around these initially sampled points. While these methods can offer sufficient diversity in solutions, they fail to fully utilize the historical information and often require a large number of evaluations for sampled configuration points to find the optimal solution. However, for tuning the inference engine, evaluations involve stress-testing the service, leading to a high time overhead. Therefore, random sampling and meta-heuristics lack efficiency for tuning the inference engine.
II-D2 Heuristic Search Algorithms
Heuristic search algorithms always leverage rules to guide the tuning process, where the search rules are often designed according to professional domain knowledge and elaborate pre-profiling to model the effects of each parameter on performance metrics. However, with the repaid development of LLM inference techniques, inference engines update frequently. For example, vLLM updates almost every few days or weeks, which leads to updates in the number and ranges of parameters, as well as changes in the effect of parameters on performance metrics. Therefore, the rules are outdated frequently, requiring substantial manpower to develop new rules. Consequently, heuristic algorithms are not suitable for tuning the LLM inference engine.
II-D3 Existing Learning-based Approaches
RL [12] and BO [13] can effectively utilize historical information to intelligently find optimal parameter configurations without prior knowledge and additional manpower. However, RL requires a time-consuming training process to obtain a well-behaved tuning agent while existing BO solutions fail to address the three unique challenges mentioned before.
TABLE III: NOTATIONS OF PROBLEM FORMULATION.
Notation Description
$\boldsymbol{\Lambda}$ The search space of parameter configurations
$\boldsymbol{x}$ A vector denotes a specific parameter configuration
$\boldsymbol{x}[i]$ The specific value for the $i$th parameter of $\boldsymbol{x}$.
$T(\boldsymbol{x})$ The objective function of request throughput for $\boldsymbol{x}$
$L(\boldsymbol{x})$ The objective function of tail request latency for $\boldsymbol{x}$
$\Phi(\boldsymbol{x})$ The objective function of average TTFT for $\boldsymbol{x}$
$\Theta(\boldsymbol{x})$ The objective function of average TPOT for $\boldsymbol{x}$
$\lambda_{t},\lambda_{l},\lambda_{\theta},\lambda_{\phi}$ Binary variables to control optimization objectives
$\mathcal{C}$ The set of known constraints
$POF(\boldsymbol{x})$ The probability of feasibility (POF) for $\boldsymbol{x}$
$\Delta$ POF threshold to mitigate hidden constraint violations
III Problem Formulation
Given an inference service, the goal of performance tuning for the LLM inference engine is to find the optimal or near-optimal parameter configuration that maximizes the objective and satisfies known and hidden constraints. As we mentioned before, different customers want to optimize different performance metrics. Therefore, we provide a generalized problem formulation that supports various optimization objectives and complex constraints. The main notations utilized in the formulation are listed in Table III.
Figure 3: SCOOT workflow. SCOOT leverages BO to find optimized parameter configurations via exploration and exploitation.
We define the configuration space $\boldsymbol{\Lambda}=\Lambda_{1}\times\Lambda_{2}\times...\times\Lambda_{n}$, where $\Lambda_{i}$ represent the range of the $i$th parameter to tune. Vector $\boldsymbol{x}\in\boldsymbol{\Lambda}$ is utilized to denote a specific parameter configuration, where the $i$th element $\boldsymbol{x}[i]\in\Lambda_{i}$ corresponds to the value of the $i$th parameter. Moreover, we respectively use $T(\boldsymbol{x})$, $L(\boldsymbol{x})$, $\Phi(\boldsymbol{x})$, and $\Theta(\boldsymbol{x})$ to represent the functions of request throughput, tail latency, average TTFT, and average TPOT for the inference service under configuration $\boldsymbol{x}$. We leverage $\mathcal{C}=\{c_{1},c_{2},...,c_{m}\}$ to symbolize the set of known constraints, where $c_{i}$ is the $i$th known constraint which restricts the relationship between elements of $\boldsymbol{x}$. For example, a known constraint can be expressed as “$\boldsymbol{x}[1]<\boldsymbol{x}[3]\ \text{if}\ \boldsymbol{x}[4]\ \text{is}\ \text{False}$”. Moreover, we utilize $POF(\boldsymbol{x})$ to represent $\boldsymbol{x}$’s probability of feasibility (POF) that $\boldsymbol{x}$ won’t violate hidden constraints. Therefore, the tuning problem can be formulated as
$\mathbb{P}:\max\limits_{\boldsymbol{x}\in\boldsymbol{\Lambda}}\ \ \lambda_{t}\cdot T(\boldsymbol{x}),\lambda_{l}\cdot L(\boldsymbol{x}),\lambda_{\phi}\cdot\Phi(\boldsymbol{x}),\lambda_{\theta}\cdot\Theta(\boldsymbol{x})$ (1)
s. t.
$c_{i},\forall c_{i}\in\mathcal{C},$ (2)
$POF(\boldsymbol{x})\geq\Delta,$ (3)
where Eq. (2) and Eq. (3) respectively denotes the known and hidden constraints. $\Delta$ is the POF threshold used to avoid violations of hidden constraints. $\lambda_{t}\in\{0,1\}$ and $\lambda_{l},\lambda_{\phi},\lambda{\theta}\in\{0,-1\}$ are utilized to control the optimization objectives. If $(\lambda_{t},\lambda_{l},\lambda_{\phi},\lambda_{\theta})=(1,0,0,0)$, the optimization objective is to only maximize the request throughput, which is applicable to periodically invoked applications such as offline recommendation. If $(\lambda_{t},\lambda_{l},\lambda_{\phi},\lambda_{\theta})=(0,-1,0,0)$, the optimization objective is to only minimize the tail latency, which is appropriate for non-interactive online applications such as classification. If $(\lambda_{t},\lambda_{l},\lambda_{\phi},\lambda_{\theta})=(0,0,-1,-1)$, the optimization objective is to simultaneously minimize the TTFT and TPOT, which is suitable for interactive applications.
IV SCOOT: Solution Description
In this section, we first introduce the overall workflow of SCOOT. Then, we elaborate on how SCOOT employs SOBO and MOBO to address the tuning problem $\mathbb{P}$. Next, we introduce two novel features of SCOOT, the parallel suggestion that can improve the tuning efficiency and random forest-based POF learning that is proposed to learn hidden constraints during the tuning process. Lastly, we describe how SCOOT ensures the stability of optimized SLOs.
IV-A SCOOT Workflow
The workflow of SCOOT is depicted in Fig. 3, involving nine key steps to find optimized parameter configurations. Customers define their optimization objectives and SCOOT ❶ gathers their request traces that include both input text of requests and output text generated by the LLM. Then, SCOOT leverages Sobol sequence-based Quasi-Monte Carlo [19] to uniformly ❷ sample configurations across the search space, where the number of samples matches the search space’s dimensionality. Then, SCOOT runs the inference engine with each sampled configuration and stress tests the inference service to obtain initial observations. These observations of configuration-performance pairs are leveraged to ❸ build a random forest and ❹ construct a surrogate model to learn the $POF(\cdot)$ and predict the probability distribution of each optimization objective, respectively. Subsequently, acquisition functions are exploited to ❺ assess configurations according to the predicted results. Based on the assessment, a solver ❻ suggests multiple configurations in parallel while adhering to the known constraints and ❼ ensuring hidden constraints using $POF(\cdot)$ learned by the random forest. Lastly, the inference engine ❽ is started with each suggested parameter configuration, and the stress testing is conducted to ❾ obtain new observations to refine the random forest and the surrogate model. Steps ❸$\sim$ run iteratively, and the tuning process stops until the number of observations reaches a given threshold.
In the above workflow, the LLM, GPU number, and GPU type used for tuning are the same as the LLM inference service owned by the customer.
IV-B Bayesian Optimization-based Solution
BO is a theoretically grounded method for solving the problem of maximizing black-box objective functions. BO leverages a probabilistic surrogate model to approximate the objective function and use acquisition functions to access configuration points for suggestion. BO can avoid excessive evaluations by exploring the complex multi-dimensional search space efficiently and intelligently [13], which is suitable for solving problems with expensive evaluation overhead.
SCOOT leverages SOBO to minimize tail latency for non-interactive online applications and maximize throughput for periodically invoked applications, respectively. For interactive applications, SCOOT exploits MOBO to minimize TTFT and TPOT simultaneously. MOBO aims to find a set of parameter configurations representing the Pareto frontier that denotes the optimal tradeoffs between TTFT and TPOT.
(a) TTFT-only optimization.
(b) TPOT-only optimization.
Figure 4: Contradictory TTFT and TOPT. Decreased TTFT often leads to increased TPOT, and vice versa. The red line indicates the objective value under the default parameter configuration.
We do not linearly combine TTFT and TPOT as a single optimization objective and solve it with SOBO for the following three reasons. First, the scale of TTFT and TPOT are significantly different. TTFT is hundreds or even thousands of times larger than TPOT. Second, it is difficult to assign weights to TTFT and TPOT to make a good trade-off. Third, under stress testing, TTFT and TPOT are always two conflicting optimization objectives.
We separately take TTFT and TPOT as the only optimization objectives and conduct performance tuning for inference services of LLAMA2-13B deployed on various computing resources using the BOT request trace. As shown in Fig. 4, the experimental results confirm that for TTFT and TPOT, optimizing for one always compromised the other. For request traces of other applications, we also observe similar results.
The reason is that in order to decrease TPOT, it is significant to decrease the batch size controlled by the parameter max-num-seqs to reduce the per-token latency, which greatly hurts the request throughput and causes a long queuing time for requests, thus leading to a large TTFT. Conversely, to reduce TTFT, a large batch size is always set to allow newly arrived requests to be processed promptly without queuing for a long time. However, a large batch size often slows down inference, leading to increased TPOT.
Therefore, for interactive online applications, optimizing the Pareto frontier of TPOT and TTFT is appropriate. Besides, this can offer customers a wide range of configurations to choose from, providing a high flexibility of SLO optimization.
IV-C Surrogate Model
The surrogate model of BO is responsible for predicting the objective function $f(\boldsymbol{x})$ on the basis of observations. It models $f(\boldsymbol{x})$ for a given configuration $\boldsymbol{x}$ as a random variable and predicts its probability distribution. In the context of performance tuning for LLM inference engines, $f(\boldsymbol{x})$ can represent the objective functions of request throughput $T(\boldsymbol{x})$, request tail latency $L(\boldsymbol{x})$, average TTFT $\Phi(\boldsymbol{x})$, and average TPOT $\Theta(\boldsymbol{x})$.
SCOOT uses the Gaussian process (GP) as the surrogate model because it is parameter-free and provides closed-form calculation. GP is characterized by a mean function $m(\boldsymbol{x})$ and a covariance function $k(\boldsymbol{x},\boldsymbol{x^{\prime}})$. Formally, a GP is defined as
$f(\boldsymbol{x})\sim GP(m(\boldsymbol{x}),k(\boldsymbol{x},\boldsymbol{x^{\prime}})),$ (4)
where the mean function $m(\boldsymbol{x})$ corresponds to the expected value of $f(\boldsymbol{x})$ for a specific input $\boldsymbol{x}$, and the covariance function $k(\boldsymbol{x},\boldsymbol{x^{\prime}})$ models the similarity between objective values at different inputs $\boldsymbol{x}$ and $\boldsymbol{x^{\prime}}$. $k(\boldsymbol{x},\boldsymbol{x^{\prime}})$ can capture the correlation structure of the objective function and control the variability and smoothness of the GP. SCOOT employs the Matern 3/2 kernel function with input wrapping as the covariance function due to its powerful capability to balance the smoothness and flexibility when modeling unknown functions [20, 21].
Given a parameter configuration $\boldsymbol{x}$, GP assumes that the probability distribution of $f(\boldsymbol{x})$ follows a Gaussian distribution whose mean $\mu(f(\boldsymbol{x}))$ is calculated by
$\mu(f(\boldsymbol{x}))=m(\boldsymbol{x})+k(\boldsymbol{x},\boldsymbol{\mathcal{X}})(k(\boldsymbol{\mathcal{X}},\boldsymbol{\mathcal{X}})+\tau^{2}I)^{-1}(\boldsymbol{\mathcal{Y}}-m(\boldsymbol{\mathcal{X}})),$ (5)
and the variance $\sigma^{2}(f(x))$ is computed by
$\sigma^{2}(f(\boldsymbol{x}))=k(\boldsymbol{x},\boldsymbol{x})-k(\boldsymbol{x},\boldsymbol{\mathcal{X}})(k(\boldsymbol{\mathcal{X}},\boldsymbol{\mathcal{X}})+\tau^{2}I)^{-1}k(\boldsymbol{\mathcal{X}},\boldsymbol{x}),$ (6)
where $\tau^{2}$ is the level of white noise, $\boldsymbol{\mathcal{X}}$ represents the previously evaluated, configurations and $\boldsymbol{\mathcal{Y}}$ denotes the corresponding observed objective values.
For SOBO, a single-output GP is leveraged to predict the objective function. For MOBO, SCOOT adopts a multi-output GP by considering each output to be independent. During the tuning process, the prediction accuracy of the GP model is continuously improved as more observations are collected.
IV-D Acquisition Function
The acquisition function of BO assesses parameter configurations in the search space based on the predicted results provided by the surrogate model. The acquisition function calculates a score for a configuration point $\boldsymbol{x}$ according to the GP’s predicted mean $\mu(f(\boldsymbol{x}))$ (indicating the expected performance) and predicted variance $\sigma^{2}(f(\boldsymbol{x}))$ (representing uncertainty). The parameter configurations with high scores are more likely to be suggested for the next evaluation. Different acquisition functions balance exploration (visiting areas with high uncertainty) and exploitation (intensively searching areas with good known objective values) in different manners.
IV-D1 SOBO Acquisition Function
For SOBO, commonly used acquisition functions include upper confidence bound (UCB), probability of improvement (PI), and expected improvement (EI). UCB incorporates both mean and variance into the score and prioritizes the point with a high balance of exploration and exploitation through a tunable trade-off parameter $\beta$, which is expressed by
$UCB(\boldsymbol{x})=\mu(f(\boldsymbol{x}))+\beta\cdot\sigma(f(\boldsymbol{x})),$ (7)
where $\sigma(f(\boldsymbol{x}))$ represents the standard deviation and it is calculated by $\sqrt{\sigma^{2}(f(\boldsymbol{x}))}$.
PI prioritizes the configuration point that is likely to yield an improvement over the current known best observation, which is expressed by
$PI(\boldsymbol{x})=P(f(\boldsymbol{x})+\xi>f(\boldsymbol{x}^{+})),$ (8)
where $\boldsymbol{x}^{+}$ is the configuration point that has the largest objective function value known so far, and $\xi$ is a parameter leveraged to encourage exploration.
EI not only considers the probability of objective improvement but also the magnitude of the expected improvement, which is computed by
$EI(\boldsymbol{x})=\mathbb{E}(\max(0,f(\boldsymbol{x})+\xi-f(\boldsymbol{x}^{+}))),$ (9)
where $\boldsymbol{x}^{+}$ is the best observed configuration point, and $\xi$ is also used to encourage exploration.
(a) Hypervolume (HV).
(b) HV improvement.
Figure 5: An illustration of two-dimensional HV. $f_{1}$ and $f_{2}$ are two objective functions and $\boldsymbol{r}$ is the reference point. (a) The blue area represents the HV of the existing solution set. (b) The yellow area depicts the HV improvement after adding $\boldsymbol{y}_{4}$.
IV-D2 MOBO Acquisition Function
For MOBO, expected hypervolume improvement (EHVI) is widely adopted as the acquisition function to identify promising parameter configurations that balance the conflicting objectives. EHVI assesses a new configuration’s quality based on the potential improvement it might bring to the hypervolume of the existing solution set, and EHVI is calculated by
$EHVI(\boldsymbol{x})=\mathbb{E}(\max(0,HV(\boldsymbol{\mathcal{Y}}\cup\{f(\boldsymbol{x})\})-HV(\boldsymbol{\mathcal{Y}}))),$ (10)
where $HV(\cdot)$ is the hypervolume function and $\boldsymbol{\mathcal{Y}}$ denotes the objective values of previously evaluated configurations.
The hypervolume function $HV(\cdot)$ is calculated by measuring the volume of space enclosed between the Pareto frontier and a reference point $\boldsymbol{r}$, where $\boldsymbol{r}$ is typically a lower bound point for all objective functions. When there are only two optimization objectives, $HV(\cdot)$ is calculated by summing the areas of rectangles formed between each point on the Pareto frontier and the selected reference point $\boldsymbol{r}$. Figure 5 presents how the two-dimensional $HV(\cdot)$ is computed.
A larger $HV(\cdot)$ value indicates a superior Pareto front as it covers a larger objective space, offering a broader range of promising solutions. Therefore, by maximizing the EVHI, the search process can be guided towards configurations that can improve the overall quality and diversity of the Pareto frontier, thus optimizing multiple objectives simultaneously and providing more excellent optional solutions.
IV-E Configuration Suggestion
IV-E1 SOBO Suggestion
For SOBO, a common practice is to select a specific acquisition function and suggest parameter configurations that maximize the acquisition function using a solver. However, when it comes to real-world applications, the selected acquisition function might work sub-optimally for the tuning task, and the best-performing acquisition function is challenging to identify in advance [22].
To mitigate this issue, SCOOT employs multi-objective acquisition function ensemble (MACE) [23]. MACE uses three acquisition functions, UCB, PI, and EI, to assess configuration points and runs a solver to find out the Pareto frontier that denotes the optimal trade-offs of the three acquisition functions. Suggested configurations are randomly selected in the Pareto frontier. With MACE, SCOOT can avoid using sub-optimal acquisition functions all the time, thereby improving the quality of suggested parameter configurations.
During the tuning process, for UCB, $\beta$ is dynamically adjusted, and in the $t$th tuning iteration, $\beta$ is calculated by ${2log(\frac{t^{2}\pi^{2}}{6\delta})}$, where $\delta=\frac{1}{t^{2}}$ [24]. For EI and PI, $\xi$ is set to a fixed value of 0.0001.
IV-E2 MOBO Suggestion
For MOBO, SCOOT leverages the EHVI as the only acquisition function and runs a solver to suggest parameter configurations by maximizing EHVI, where the default parameter configuration of the inference engine is chosen as the reference point $\boldsymbol{r}$.
IV-E3 Known Constraint Assurance
For both SOBO and MOBO, in the process of solvers solving optimization problems, the solution space is pruned by known constraints to avoid engine startup failure, and hence the known constraints can be guaranteed.
Input: $N$: the total number of observations;
$k$: parallelism degree;
1
// Initialize observations
2 $\mathcal{O}\leftarrow$ UNIFORM_SAMPLE_AND_EVALUATE()
// Train random forest regression
$RF\leftarrow$ TRAIN($\mathcal{O}$) // Train Random Forest
3
[…]
5while $|\mathcal{O}|<N$ do
       // Suggest $k$ configurations
6       $\{\boldsymbol{x_{1}},...,\boldsymbol{x_{k}}\}\leftarrow$ SUGGEST($k$, $\mathcal{O}$, $RF$, $\Delta$)
      // Collect performance metrics
7       $\{\boldsymbol{y_{1}},...,\boldsymbol{y_{k}}\}\leftarrow$ EVALUATE($\{\boldsymbol{x_{1}},...,\boldsymbol{x_{k}}\}$)
      // Adjust feasibility threshold
8       if there are invalid $\boldsymbol{y}$ in $\{\boldsymbol{y_{1}},...,\boldsymbol{y_{k}}\}$ then
9             $\Delta\leftarrow\min(0.75,\max(0.5,\Delta+\upsilon))$
            $c$ $\leftarrow 0$ /* Clear continuous feasible suggestion count */
10
11       else
12             $c\leftarrow c+k$
13            if $c\geq 5$ then
14                   $\Delta\leftarrow\max(0.25,\Delta-\upsilon)$
15                  $c\leftarrow c-5$
16
17
      // Update observations
18       $\mathcal{O}\leftarrow$ UNION($\mathcal{O}$, $\{\boldsymbol{x_{1}},...,\boldsymbol{x_{k}}\}$, $\{\boldsymbol{y_{1}},...,\boldsymbol{y_{k}}\}$)
      $RF\leftarrow$ TRAIN($\mathcal{O}$) // Refine RF
19
Algorithm 1 SCOOT: BO-based Performance Tuning
IV-F Parallel Configuration Suggestion
With the popularity of LLMs, an increasing number of LLM inference services are being deployed, and tuning them incurs substantial time overhead. To mitigate this problem, SCOOT suggests multiple parameter configurations at a time and utilizes multiple sets of computing resources to conduct configuration evaluations in parallel.
Given the parallelism degree $k$, for SOBO, SCOOT randomly selects $k$ points from the Pareto frontier of the three acquisition functions for suggesting. For MOBO, the top-$k$ points with the largest EHVI are chosen for suggesting. Thus, tuning can be accelerated by a factor of $k$ when the number of observations is fixed.
Although parallel suggestions require additional computing resources to evaluate configurations in parallel, many computing resources in production clusters are often idle during off-peak hours such as nights and weekends, which can be conveniently used for tuning the inference engine.
IV-G Random Forest-based POF Learning
Frequent violations of hidden constraints can result in a significant number of invalid observations caused by engine crashes, which not only hurts the tuning efficiency but also hampers the surrogate model from adequately learning the relationship between the objective functions and configuration points, thereby hurting the optimization of SLOs. To handle hidden constraints, a common practice is to assign penalty objective values to the parameter configurations that cause engine crashes. However, it is challenging to set appropriate penalty objectives and such penalty objective values often negatively affect the surrogate model [25]. Therefore, SCOOT leverages random forest regression to learn the $POF(\cdot)$ for parameter configurations and handle hidden constraints by limiting the probability of feasibility to be greater than or equal to a given threshold, which is expressed by Eq. (3) in Section III. By restricting the solver from adhering to the hidden constraints during the configuration suggestion, SCOOT can substantially reduce the number of invalid observations without affecting the surrogate model.
For the hidden constraint expressed in Eq. (3), the feasibility probability threshold $\Delta$ is critical. If $\Delta$ is fixed, setting it too high may result in repeatedly suggesting configuration points near feasible configurations, potentially missing out on other superior configuration points. Conversely, setting the threshold too low may lead to suggestions of many infeasible configurations. Therefore, SCOOT dynamically adjusts the threshold $\Delta$ during the tuning process.
The dynamic adjustment process of $\Delta$ can be found in Algorithm 1, where we can observe that when infeasible parameter configurations are suggested, $\Delta$ is increased to reduce the likelihood of further suggesting infeasible configurations. Besides, after continuously suggesting feasible configurations five times, $\Delta$ is decreased to allow the discovery of potential superior configurations. Additionally, during the tuning process, the random forest is continuously refined using the latest observations to enhance prediction accuracy, ensuring that hidden constraints are intelligently followed without compromising the quality of performance tuning.
(a) BOT.
(b) SQL.
(c) CLS.
(d) REC.
Figure 6: Probability density function (PDF) of the request input and output lengths for the four request traces used in experiments.
Since known and hidden constraints pure the solution space, solvers may be unable to give a sufficient number of solutions during the configuration suggestion. In such scenarios, SCOOT randomly samples configuration points with Sobol sequence-based Quasi-Monte Carlo to make up for the shortfall.
IV-H Ensuring SLO Stability
Since the order in which requests arrive can affect the performance metrics obtained by stress testing, the SLOs of performance tuning vary with different request arrival orders. Through extensive experiments, we find that although the optimal objective value varies with the requests’ arrival order, the corresponding parameter configurations are often very similar. Consequently, after completing the tuning process, we re-conduct the stress testing multiple times with different request arrival orders and select the worst-case objective value as the SLO to ensure the stability of the SLOs. Finally, we summarize the overall tuning procedure of SCOOT in Algorithm 1.
V Experimental Evaluation
V-A Experiment Setup
Implementation. We implement SCOOT with Python 3.8. Pytorch [26] and GPytorch [27] are used to implement the Sobol sequence-based Quasi-Monte Carlo sampling and GP, respectively. For configuration suggestion, we implement the solver with pymoo [28]. Moreover, the random forest regression is implemented using sklearn [29].
Computing Resources. Experiments are conducted on our testbed. Multiple sets of computing resources are utilized, including 2A10, 4A10, 2A100, and 4A100, where A10 represents the NVIDIA A10 24GB GPU, A100 denotes the NVIDIA A100 80GB GPU, and the integer before the GPU type indicate the number of GPUs. For each set of computing resources, 256GB CPU memory and one 2.90GHz 32-core Intel(R) Xeon(R) CPU are equipped. Besides, A100 GPUs are connected over NVLink, and A10 GPUs are connected over PCIE.
Request Trace. We collect request traces from four LLM inference services at AntGroup, which include four applications of text-to-sql (SQL), chatbot (BOT), classification (CLS), and recommendation (REC). SQL and BOT are interactive applications, CLS is an online application while REC is an offline application. The distribution of request input lengths and output lengths for the four LLM-empowered applications are presented in Fig. 6.
Inference Engine. We adopt vLLM-v0.4.2 as the inference engine, whose default configuration of to-be-tuned parameters is the same as the latest version.
LLM. We conducted experiments with the LLAMA2-7B and LLAMA2-13B models [18] because the model architecture of LLAMA is representative. Besides, in practice, 7B and 13B are common parameter sizes, and fine-tuned 7B and 13B LLMs can fulfill the requirements for most applications.
Workflow. In each configuration evaluation, when the inference engine is started with the suggested configuration, we send requests from the request trace for 100 seconds and the request arrival times are generated using Poisson distribution with various request rates. For different request traces and GPU types, the request rate is different. For A10, the request rate for SQL, BOT, CLS, and REC are 5,5, 10, and 15, respectively. For A100, the request rates of these request traces are twice that of A10. The request rate is set according to the actual workload of the applications as well as considering the computing capability and GPU memory capacity of GPUs.
Metrics. Since different applications aim to optimize different performance metrics. For SQL and BOT, TTFT and TPOT are leveraged as optimization objectives, while for CLS and REC, 95th percentile latency and request throughput are utilized as performance metrics, respectively. In the experimental results, we report the improvement or reduction of these metrics over the default configuration.
(a) TTFT reduction.
(b) TPOT reduction.
(c) Avg TTFT and TPOT reduction.
Figure 7: SLO optimization for BOT under LLAMA2-7B.
(a) TTFT reduction.
(b) TPOT reduction.
(c) Avg TTFT and TPOT reduction.
Figure 8: SLO optimization for BOT under LLAMA2-13B.
(a) TTFT reduction.
(b) TPOT reduction.
(c) Avg TTFT and TPOT reduction.
Figure 9: SLO optimization for SQL under LLAMA2-7B.
(a) TTFT reduction.
(b) TPOT reduction.
(c) Avg TTFT and TPOT reduction.
Figure 10: SLO optimization for SQL under LLAMA2-13B.
Baselines. We compare SCOOT with the following baselines.
  • Random Sampling (RD): Sobol sequence-based Quasi-Monte Carlo [19] is utilized to uniformly sample configuration points from the search space.
  • Genetic Algorithm (GA): Mixed variable GA [30] is leveraged to reduce request tail latency and improve request throughput for CLS and REC applications, while NSGA2 [31] can handle multi-objective problems and is utilized to optimize TTFT and TPOT simultaneously for BOT and SQL applications. For GA, we assign a penalty objective value when constraints are violated. Besides, the population size of GA is set to ten.
  • Heteroscedastic Evolutionary BO (HEBO): HEBO [22] is a leading performance tuning framework that exploits Bayesian optimization to solve black-box optimization problems. HEBO can handle single-objective and multi-objective optimization problems with SOBO and MOBO, respectively. For HEBO, we assign a penalty objective value when constraints are violated.
(a) Tail latency reduction under LLAMA2-7B.
(b) Tail latency reduction under LLAMA2-13B.
(c) Avg tail latency reduction.
Figure 11: SLO optimization for CLS.
(a) Request TP improvement for LLAMA2-7B.
(b) Request TP improvement for LLAMA2-13B.
(c) Avg request TP improvement.
Figure 12: SLO optimization for REC.
V-B SLO Optimization
In the experiments of SLO optimization, we limit the total number of observations to 30 and set the suggestion parallelism degree (PD) for SCOOT to one. The performance of SLO optimization for SCOOT and baselines is presented in Figures 7, 8, 9, 10, 11, and 12. The bar represents the performance gain obtained by tuning, and the higher the bars, the better. In addition, × represents that no better configuration than the default configuration is found during the tuning process.
For BOT and SQL applications, TTFT and TPOT are optimized simultaneously. We present the tuning results of the two applications under different LLMs and computing resources, as well as the average tuning results across various computing resources in Figures 6, 7, 8, and 9. Parameter configurations obtained by tuning form a Pareto frontier of TTFT and TPOT, and we choose the configuration with the largest HV from the Pareto frontier to report since it represents the best trade-off between TTFT and TPOT. HV of a single configuration point $\boldsymbol{x}$ is calculated by
$HV(\boldsymbol{x})=\max(0,y_{\boldsymbol{x}}^{\phi}-y_{\boldsymbol{r}}^{\phi})\cdot\max(0,y_{\boldsymbol{x}}^{\theta}-y_{\boldsymbol{r}}^{\theta}),$ (11)
where $y_{\boldsymbol{x}}^{\phi}$, $y_{\boldsymbol{x}}^{\theta}$, $y_{\boldsymbol{r}}^{\phi}$, and $y_{\boldsymbol{r}}^{\theta}$ denote the observed TTFT and TPOT of $\boldsymbol{x}$ and the reference point $\boldsymbol{r}$ (i.e., the default configuration of the LLM inference engine), respectively.
From Fig 7, Fig. 8, Fig. 9, and Fig. 10, we can observe that for the average tuning performance in SLO optimization, SCOOT surpasses all the baselines in terms of both TTFT reduction and TPOT reduction under different LLMs, which confirms the SCOOT’s superiority in SLO optimization for interactive applications. For the BOT application, SCOOT decreases the TTFT and TPOT by up to 98.9% and 10.5% compared to the default configuration and baselines, respectively. In addition, for the SQL application, SCOOT reduces the TTFT and TPOT by up to 99.8% and 61.0% compared to the default configuration and baselines, separately.
While in some situations, SCOOT only outperforms the baselines on either TTFT or TPOT reduction, the optimized SLOs of SCOOT have a larger HV. For example, as shown in Fig. 8, in the case of 2A100, HEBO reduces TPOT slightly more than SCOOT, but SCOOT significantly outperforms HEBO in TTFT reduction, which leads to the optimized SLOs of SCOOT having an HV that is considerably larger than that of HEBO. Additionally, although in a very small number of cases, SCOOT failed to outperform the baseline in both TTFT and TPOT reductions, SCOOT’s tuning performance is more consistent. For example, as shown in Fig. 10, in the case of 2A100, HOBO outperforms SCOOT in both TTFT and TPOT reductions. However, in the case of 4A10, HEBO fails to find a configuration that is better than the default configuration, while SCOOT greatly reduces both TTFT and TPOT. In conclusion, SCOOT can always find superior parameter configurations. Such a consistently superior performance makes SCOOT outperform all the baselines in the average performance of SLO optimization across various computing resources.
For CLS and REC applications, request tail latency and request throughput are respectively utilized as optimization objectives. We present the tuning results of the two applications under different LLMs and computing resources in Figures 11 and 12, where we can observe that SCOOT always surpasses baselines in terms of tail latency reduction and throughput improvement under different LLMs, which confirms the SCOOT’s superiority in single-objective optimization.
For the CLS application, SCOOT reduces request tail latency by up to 40.6% compared to the default configuration and baselines. For the REC application, SCOOT decreases request throughput by up to 68.3% compared to the default configuration and baselines. Although HEBO’s performance is comparable to SCOOT and even slightly exceeds SCOOT in the case of 2A100 under LLAMA2-7B, as shown in Fig. 12, for LLAMA2-13B, HEBO’s performance is always far inferior to SCOOT, and hence SCOOT outperforms HEBO in average performance of SLO optimization for the REC application. This also confirms the SCOOT’s consistently superior performance on single-objective optimizations.
To sum up, for SLO optimization, SCOOT is always capable of finding superior parameter configurations for LLM inference services of various requests, LLMs, and computing resources to optimize various performance metrics.
(a) SLO optimization for BOT.
(b) SLO optimization for CLS.
Figure 13: Effectiveness of random forest-based POF learning.
SCOOT’s outstanding performance primarily comes from two aspects. First, SCOOT employs BO, which can efficiently handle black-box optimization problems with expensive evaluation overhead. Second, SCOOT can tackle complex constraints. By using hard constraints and learned hidden constraints to prune the solver’s solution space, it can significantly reduce the number of invalid observations without affecting the surrogate model. Additionally, this allows the surrogate model to fully learn the relationship between parameter configurations and optimization objectives. Therefore, SCOOT can always find good configurations within 30 observations.
GA optimizes around the initially sampled configurations. With a limited number of observations, both the population size and the number of iterations of GA are restricted to a very small value. Therefore, GA not only lacks diversity in solutions but also cannot fully utilize historical information to guide the search process, hence often performing poorly.
RD is unable to utilize historical information to guide the sampling process and can not handle constraints. Therefore, RD often fails to obtain sufficiently good configurations within 30 observations. However, due to the uniform sampling, RD can explore a large range in the search space and hence has a higher diversity in sampled configurations than GA. Hence, RD can often perform better than GA.
By leveraging BO to intelligently utilize historical data to guide the tuning process, HEBO often outperforms RD and GA in SLO optimization. However, HEBO simultaneously learns the objective function and constraints using the surrogate model. By assigning a penalty objective value to the unfeasible configurations, the learning of constraints hurts the modeling of the objective function. Configuration points near infeasible configurations are considered to be poorly performing points and are confused with values that are feasible but poorly performing. This can lead to a lack of subsequent exploration in the adjacent area of infeasible configurations, resulting in some potentially superior solutions being missed.
SCOOT can address this issue by leveraging GP as the surrogate model to approximate objective functions while learning hidden constraints with random forest, and hence the learning of hidden constraints does not affect the modeling of objective functions. In addition, by dynamically adjusting the POF threshold $\Delta$, SCOOT can explore the region near the infeasible configurations after continuous exploration of feasible configurations so as to explore the solution space more fully. We conduct experiments to confirm the effectiveness of the random forest-based learning of hidden constraints under 2A100 and 4A100. The results are presented in Fig. 13. Since requests from the CLS application seldom cause engine crashes, not applying RF-based POF learning has little impact on SCOOT’s performance. However, since BOT application requests cause many hidden constraints, RF-based POF learning can greatly enhance SLO optimization.
V-C Solution Quality
For multi-objective optimization, we assess the quality of solutions from two perspectives. First, the optimality of solutions, whether optimal trade-offs have been achieved between TTFT and TPOT, that is, whether the optimized parameter configurations lie on the Pareto frontier of TTFT and TPOT. Second, we focus on the diversity of solutions, that is, whether the range of optimized parameter configurations on the Pareto frontier is sufficiently broad to offer customers more optional promising configurations. A high diversity of solutions can enhance the flexibility of SLO optimization.
In Fig. 14 and Fig. 15, we plot the Pareto sets of configuration points sampled in the process of tuning the BOT and SQL applications under 4A10 for SCOOT and baselines. The plotted points represent the “local optimal solutions” identified by each tuning method. Additionally, we mark the global Pareto set from these local optimal configuration points with $\boldsymbol{\star}$. From Fig. 14, we can observe that for BOT, the majority of global optimal configuration points are contributed by SCOOT, both under LLAMA-13B and LLAMA-7B. Additionally, SCOOT yields a wider range of these global optimal points compared with baselines. From Fig. 15, we can see that for SQL, all the global optimal configuration points are contributed by SCOOT under both LLAMA-13B and LLAMA-7B. These experimental results confirm the superiority of SCOOT in both the optimality and diversity of SLO optimization in multi-objective situations.
(a) Configs under LLAMA2-7B.
(b) Configs under LLAMA2-13B.
Figure 14: Solution quality for BOT.
(a) Configs under LLAMA2-7B.
(b) Configs under LLAMA2-13B.
Figure 15: Solution quality for SQL.
(a) SLO optimization for CLS.
(b) SLO optimization for BOT.
(c) Performance tuning time for BOT and SQL.
Figure 16: Tuning efficiency of various parallelism degrees (PDs).
V-D Tuning Efficiency
In this subsection, we evaluate the tuning efficiency for SCOOT from two aspects. First, we focus on the speedup brought by the parallel suggestion. Second, we evaluate the efficiency of BO algorithms in making suggestions. We conduct experiments to tune the CSL and BOT applications under 2A100 with different parallelism degrees. In the experiments, the number of observations is also limited to 30 when the parallelism degree is set to 2. Figures 16(a) and 16(b) show the SCOOT’s performance in SLO optimization under various parallelism degrees, while Fig. 16(c) presents the total tuning time across various parallelism degrees.
From Figures 16(a) and 16(b), we can observe that the parallel suggestion achieves the same performance in SLO optimization as the non-parallel suggestion when the parallelism degree is set to 2. Additionally, as seen in Fig. 16(c), when setting the parallelism degree to 2, the parallel suggestion technique significantly reduces the total tuning time, nearly cutting it in half, which demonstrates that parallel suggestions can effectively improve tuning efficiency. Furthermore, from Fig. 16(c), we can also find that compared with the time of configuration evaluation, the time required for configuration suggestion is negligible, which validates the efficiency of applying BO algorithms to tune LLM inference engines. Moreover, the time for configuration suggestion also drastically decreases with the parallel suggestion. This is because, for both multi-objective and single-objective optimization, the parallel suggestion technique only samples additional points in the final sampling step, incurring negligible computational overhead. Since the number of recommended points is doubled, with the total number of observations fixed, the number of configuration suggestions is reduced to half, resulting in the time required for suggesting configuration points being cut in half as well. Compared to baselines, SCOOT can utilize more computational resources to accelerate the tuning process using the parallel suggestion, offering greater flexibility in efficiency improvement.
We also increase the parallelism degree and conduct experiments. However, we find that when the parallelism degree is set to a large value, such as greater than 4, SCOOT’s performance of SLO optimization is compromised. This is because when the parallelism degree increases, the surrogate model in each configuration suggestion is not as accurate as when the parallelism degree (PD) is set small. Let’s consider the following scenario. During the $i$th tuning iteration, the size of the observation set is $n$. For a case where PD=$k(k>1)$, SCOOT suggests $k$ configurations in one go, denoted as the $n+1$th, $n+2$th,…, $n+k$th configuration. All the configurations are suggested based on the surrogate model built with $n$ observations. When PD is small, for example, PD=1, the $n+1$th configuration is suggested based on the previous $n$ observations, but the $n+2$th configuration is suggested based on $n+1$ observations, making the prediction more accurate.
Since the parallel suggestion can reduce the total tuning time by PD times when the number of observations is fixed, a relatively small PD setting that does not affect the optimization results can significantly improve tuning efficiency. In addition, we also experimentally find that when the PD is set to a large value, the problem of performance degradation in SLO optimization can be resolved by appropriately increasing the number of observations. In the future, we will explore the best number of observations under different parallelism degrees to find the most cost-effective acceleration solution.
VI Related Works
VI-A Advanced LLM Inference Techniques
Existing inference engines such as vLLM and TensorRT-LLM support a variety of advanced inference techniques to improve the speed and throughput of LLM inference, including a series of kernel-level optimizations such as continuous batching [6], paged attention [7], flash attention [32, 33], and chunked prefill [8]. Besides, prefix caching [34] is employed by inference engines to reuse computed key and value tensors of requests’ common prefix texts to reduce redundant memory allocation and computation for newly arrived requests, which can significantly improve the serving efficiency when requests share a long common prefix. Moreover, speculative decoding [17, 35, 36, 37, 38] is also adopted by inference engines to accelerate LLM inference, where an efficient draft model is leveraged to generate tokens, and the LLM occasionally refines the draft.
Some of these technologies are effective in specific scenarios and have been implemented as configurable parameters of inference engines. However, how to combine these technologies to fully exploit the capabilities of inference engines has not been adequately studied, a gap that this study aims to fill.
VI-B Automatic Performance Tuning Approaches
Performance tuning techniques have been widely applied across a wide variety of fields, including database tuning [39, 40, 41], Spark configuration tuning [42, 43, 44], and compiler optimization [25]. The vast majority of these performance tuning studies adopt BO to find optimized parameter configurations since BO is theoretically grounded and can efficiently learn the relationship between performance and parameters from evaluations, thus intelligently tuning parameters for performance improvement. Because of these advantages, BO is also employed in resource allocation studies [45, 46].
However, to the best of our knowledge, previous studies have not focused on tuning the parameters of LLM inference engines and can not address the three unique challenges faced in LLM inference engine tuning.
VII Conclusion and Future Work
In this paper, we propose SCOOT, a BO-based automatic inference engine tuning system, to optimize SLOs for LLM inference services. SCOOT leverages both SOBO and MOBO to handle various optimization objectives and utilizes random forest regression to learn hidden constraints during the tuning process. Additionally, SCOOT exploits the parallel suggestion to accelerate the tuning process using additional computing resources. Experimental results show that SCOOT can effectively speed up the tuning process and significantly optimize SLOs, improving the request throughput by up to 68.3%, reducing the request tail latency by up to 40.6%, and reducing the TTFT and TPOT by up to 99.8% and 61.0%, respectively, compared to the default parameter configuration and existing tuning methods. In the future, we intend to expand SCOOT’s capabilities to accommodate more inference engines, such as TensorRT-LLM.